OpenStudy (anonymous):

The length of a rectangle is 10 more than three times the width. The perimeter is 108. Find the dimensions

OpenStudy (anonymous):

10+3w = l P = 2l +2w

OpenStudy (anonymous):

P = 2(10+3w) + 2w

OpenStudy (anonymous):

108 = 10 +6w +2w

OpenStudy (anonymous):

98 = 6w + 2w

OpenStudy (anonymous):

98 = 8w

OpenStudy (anonymous):

w = 12.25

OpenStudy (anonymous):

l = 41.75

OpenStudy (anonymous):

P = 108

OpenStudy (anonymous):

width = 12.25 length = 41.75 Perimeter = 108